Ram Corp's last dividend was $1.75 (D0). Its dividend growth rate is expected to be at 25% for each of the first 2 years, after which dividends are expected to grow at a constant rate of 6% forever. Its required rate of return is 14%. What is its stock price?

A)

$32.76

B)

$30.48

C)

$35.45

D)

$31.90

E)

$34.60
Answer #1

Solution:-

DPS0 = $1.75

DPS1 = DPS0 * (1+G)

DPS1 = $1.75 * (1+0.25)

DPS1 = $2.1875

DPS2 = DPS1 * (1+G)

DPS2 = $2.1875 * (1+0.25)

DPS2 = 2.7344

To Calculate Terminal Value at the year end 2-

Terminal Value = \frac{DPS2 * (1+G)}{Required Return - Growth Rate}

Terminal Value = \frac{2.7344 * (1+0.06)}{0.14 - 0.06}

Terminal Value = $36.231

To Calculate Stock Price-

Stock Price = \frac{DPS1}{(1+r)^{1}} + \frac{DPS2 + Terminal Value}{(1+r)^{2}}

Stock Price = \frac{2.1875}{(1+0.14)^{1}} + \frac{2.7344 + 36.231}{(1+0.14)^{2}}

Stock Price = 2.1875 * 0.877 + 38.9654 * 0.769

Stock Price = $1.91886 + $29.98

Stock Price = $31.90

The Correct Answer is point D i.e. $31.90

If you have any query related to question then feel free to ask me in a comment.Thanks. Please rate.